OpenStudy (anonymous):

Can someone verify I've done this right? Lim x->0 of (sin^2 (5x))/7x^2. I've come up with 25/49.

myininaya (myininaya):

\[\lim_{x \rightarrow 0}\frac{\sin(5x) \cdot \sin(5x)}{7 \cdot x \cdot x}=\frac{25}{7}\lim_{x \rightarrow 0}\frac{\sin(5x)}{5x} \cdot \frac{\sin(5x)}{5x} \]

OpenStudy (anonymous):

the 49 is a mistake

myininaya (myininaya):

yes it is just 7 on bottom

OpenStudy (anonymous):

what the talented myininaya said

OpenStudy (anonymous):

Devilry! I accidently multiplied 7x twice. I shouldn't be trusted with algebra. Thanks.

myininaya (myininaya):

no one should be trusted with algebra

myininaya (myininaya):

its all to get us all

myininaya (myininaya):

i'm jk mistakes happen lol